OpenStudy (anonymous):

Don and Jacob want to know who reads more pages per hour. Don keeps track of the number of pages he reads in a graph, and Jacob keeps track of how many pages he reads in a table. How many pages per hour does each boy read, and who reads at a faster rate? Don: pages per hour Jacob: pages per hour The person who reads faster is: (Type D for Don or J for Jacob) Jacob’s Reading Log Time (h) 2 3.5 5 7 Pages read 42 73.5 105 147

OpenStudy (anonymous):

@dtan5457 plz help plz

OpenStudy (danjs):

The rate of reading in pages per hour, is the slope, rate of change, of the line in the graph, and in the table`

OpenStudy (danjs):

For Don...Using any 2 points on the graph, calculating the slope... Done rate of reading = (108-72)/(6-4) = ??? pages per hour

OpenStudy (danjs):

For jacob, use any two points in the table... Jacob's rate of reading = (147 - 105) /(7 - 5) = ??? pages per hour

OpenStudy (anonymous):

108-72=36 6-4=2

OpenStudy (danjs):

yes

OpenStudy (anonymous):

36/2=18

OpenStudy (danjs):

Don reads 36 pages in 2 hours, or he reads 18 pages per hour

OpenStudy (anonymous):

147-105=42 7-5=2 42/2=21 so I think Jacob reads more

OpenStudy (danjs):

yep, that is all you have to do

OpenStudy (danjs):

Jacob reads 21 pages per hour Don Reads 18 pages per hour

OpenStudy (anonymous):

thanks